# Multi-target Range, Doppler and Angle estimation in MIMO-FMCW Radar with Limited Measurements

### v1

@@ III Sparse target-parameter recovery
In this section, unlike prior studies, we consider both range and joint Doppler-angle estimation using CS-based methods in Sections III-B and III-C, respectively. As a precursor, Section III-A provides a brief overview of the standard sparse signal recovery problem and the CS framework.
@@ III-A CS sparse signal recovery
[…]
@@ III-B Range estimation
We first estimate the target ranges from measurements (3) using two different methods: DFT-based focusing and OMP. The target velocities and AOAs are then jointly estimated for each detected range in the subsequent section.
@@
$=\sum_{k=1}^{K}\widetilde{a}^{*}_{k}\exp{(j2\pi\Omega^{k}_{D}\zeta_{p})}\exp{(j2\pi\Omega^{k}_{\theta}(\alpha_{n}+\beta_{m}))}$
×$\times\frac{1}{\sqrt{N}}\sum_{t=0}^{N-1}\exp{\left(j2\pi\left(\Omega^{k}_{R}-\frac{l}{N}\right)t\right)}+W_{n,m,p}[l],$ (9)
for $10\leq l\leq N-$ 1 and $W_{n,m,p}[l]=(1/\sqrt{N})\sum_{t=0}^{N-1}w_{n,m,p}[t]\exp{(-j2\pi lt/N)}$.
@@
$|g(x|\overline{x})|\approx\begin{cases}\sqrt{M},&|x-\overline{x}|\leq\pi/M\omega\\ 0,&|x-\overline{x}|>\pi/M\omega\end{cases}.$
This approximation implies that in the focus zone $|x-\overline{x}|\leq\pi/M\omega$, the $M$ exponents are coherently integrated while the signal outside the focus zone is severely attenuated. This focusing operation was applied across pulses in [26] to reduce the joint delay-Doppler estimation problem to delay-only estimation in a pulsed-Doppler radar. In our range estimation, the sum of exponents appears naturally in the IF signal’s DFT in (9).
@@
Binary integration: So far, we have considered range estimation from a given $p$-th chirp and $(n,m)$-th virtual array channel. Similarly, the range estimates are computed independently for all $P$ chirps and $N_{T}N_{R}$ channels using the radar measurements $\{y_{n,m,p}[t]\}_{1\leq n\leq N_{T},1\leq m\leq N_{R},1\leq p\leq P}$. In binary integration, these detected ranges are then filtered for false alarms and missed detections across all chirps and channels using a majority rule, i.e., only the ranges detected in the majority of measurements are considered valid.
@@
Classical-DFT range processing also involves peak detection in the DFT of the sampled IF signal. However, in classical processing, the measurements from all chirps and array channels are processed coherently to average out the noise term. However, at low SNRs, the method fails to detect range bins where the noise is significant. Contrarily, in binary integration, a missed range in one measurement can be detected in other measurements, reducing the missed detection probability, as demonstrated in Section V-A. However, our joint Doppler-angle estimation procedure proposed later can also be trivially applied to coherently processed measurements.
@@
Consider the measurements $y_{1,1,1}[t]$ from the first chirp and the first virtual array channel. Denote $x_{k}\doteq\widetilde{a}^{*}_{k}\exp{(j2\pi\Omega^{k}_{D}\zeta_{p})}\exp{(j2\pi\Omega^{k}_{\theta}(\alpha_{n}+\beta_{m}))}$ and stack the samples $y_{1,1,1}[t]$ for $10\leq t\leq N-$ 1 in a $N\times 1$ vector $\mathbf{y}_{R}$. Now, define the $K\times 1$ vector $\widetilde{\mathbf{x}}_{R}=[x_{1},\ldots,x_{K}]^{T}$ such that (3) yields
@@
where the $N\times K$ matrix $\widetilde{\mathbf{A}}(\bm{\Omega}_{R})=[\mathbf{a}(\Omega^{1}_{R}),\ldots,\mathbf{a}(\Omega^{K}_{R})]$ with the $k$-th column $\mathbf{a}(\Omega^{k}_{R})\doteq[1,\exp{(j2\pi\Omega^{k}_{R})},\ldots,\exp{(j2\pi\Omega^{k}_{R}(N-1))}]^{T}$. Here, $\mathbf{w}_{R}$ represents the $N\times 1$ stacked noise vector. Each column $\mathbf{a}(\Omega_{R})$ of matrix $\widetilde{\mathbf{A}}$ is parameterized by the (range) beat frequency $\Omega_{R}$ and is equivalent to steering vector in the beat frequency domain. Furthermore, $\mathbf{y}_{R}$ in (12) represents full measurements in the fast-time domain sampled at frequency $f_{s}$.
@@
where the $G_{R}\times 1$ vector $\mathbf{x}_{R}$ contains the target ranges and unknown coefficients $\{x_{k}\}$. In particular, a non-zero element of $\mathbf{x}_{R}$ represents a target present at the range corresponding to the grid point. Since $K\ll G_{R}$, $\mathbf{x}_{R}$ is a sparse vector and range estimation reduces to determining $\textrm{supp}(\mathbf{x}_{R})$ given $\mathbf{y}_{R}$ and $\mathbf{A}$. To this end, we can employ any 1D-MP or BP recovery algorithms described in Section III-A. In particular, we use OMP in Section V.
@@
Range-OMP in (13) considers measurements only from one chirp and one array channel. This is in contrast to DFT-focusing, wherein all the measurements are integrated (binary or coherently) to obtain valid target ranges making them computationally expensive. Furthermore, range-OMP provides superior range resolution compared to DFT-focusing, determined by the choice of grid $\omega_{1\leq g\leq G_{R}}$ and not by parameters $\gamma$ and $T_{c}$. However, the $\textrm{supp}(\mathbf{x}_{R})$’s recovery still depends on the matrix $\mathbf{A}$ such that the recovery probability is low if the grid points are too close.
@@ III-C Joint Doppler-angle estimation
We stack the measurements $Y_{n,m,p}[l^{\prime}]$ for all $1\leq n\leq N_{T}$ and $1\leq m\leq N_{R}$ in an $N_{T}N_{R}\times 1$ vector $\mathbf{y}_{p}$ for the $p$-th chirp. Further, define the $N_{T}N_{R}\times P$ matrix $\mathbf{Y}=[\mathbf{y}_{1},\ldots,\mathbf{y}_{P}]$ and the steering vectors in the Doppler and angular frequency domains as $\mathbf{b}(\Omega_{D})$ and $\mathbf{c}(\Omega_{\theta})$, respectively, as
@@
where the unknown $G_{\theta}\times G_{D}$ matrix $\mathbf{Z}$ contains the coefficients $\{z_{k^{\prime}}\}$ as well as the target velocities and AOAs information. Particularly, a non-zero element in $\mathbf{Z}$ represents a target with its AOA and velocity equal to that of the corresponding point in the 2D grid $\{(\phi_{i_{1}},\rho_{i_{2}}):1\leq i_{1}\leq G_{\theta},1\leq i_{2}\leq G_{D}\}$. Since $K^{\prime}\ll G_{D}G_{\theta}$, the matrix $\mathbf{Z}$ is sparse, and hence, the desired velocities and AOAs can be recovered using small number of measurements $\mathbf{Y}$ given matrices $\mathbf{C}$ and $\mathbf{B}$. Note that the joint Doppler-angle estimation has now been reduced to determining $\textrm{supp}(\mathbf{Z})$. In (18), the measurement matrices $\mathbf{C}$ and $\mathbf{B}$ and hence, the recovery guarantees depend on the choice of girds $\rho_{1\leq g\leq G_{D}}$ and $\phi_{1\leq g\leq G_{\theta}}$ as well as the number and (random) locations of the transmitted chirps ($\zeta_{1\leq p\leq P}$) and antenna elements ($\alpha_{1\leq n\leq N_{T}}$ and $\beta_{1\leq m\leq N_{R}}$).
@@
which is same as (4) with $\mathbf{A}$ and $\mathbf{x}$ replaced by $\mathbf{D}$ and $\mathbf{z}$, respectively. In Section IV, we consider the $l_{1}$-minimization-based solution of (19) to derive the theoretical recovery guarantees while Section V numerically compares both 2D-OMP and vectorized CS methods.
@@ IV-A Uniform recovery
Then, for $10<\vartheta<$ 1, the coherence $\mu$ of matrix $\mathbf{D}$ in (19) satisfies the following:
@@
×$\times(1-2\vartheta\sqrt{N_{T}N_{R}}\;\mathcal{K}_{1}(2\vartheta\sqrt{N_{T}N_{R}}))^{G_{\theta}-1},$ (27)
@@
Consider the random sparse chirps and SLA radar setup with the distributions $\mathcal{P}_{\alpha}$, $\mathcal{P}_{\beta}$ and $\mathcal{P}_{p}$ and grids $\rho_{1:G_{D}}$ and $\phi_{1:G_{\theta}}$ satisfying the conditions of Theorem 1. Also, consider some $10<\epsilon<$ 1 and two positive constants $\epsilon_{1}$ and $\epsilon_{2}$ such that $\epsilon_{1}+\epsilon_{2}=\epsilon$ and
@@
$N_{T}N_{R}\geq\kappa_{2}\left(K-\frac{1}{2}\right)^{2}\left(\log\left(\frac{G_{\theta}\sqrt{\pi}}{2\epsilon_{2}}\right)+\frac{1}{2}\log\left(2\log\left(\frac{G_{\theta}\sqrt{\pi}}{2\epsilon_{2}}\right)\right)\right)^{2},$ (30)

### v2

@@ III Sparse target-parameter recovery
In this section, unlike prior studies, we consider both range and joint Doppler-angle estimation using CS-based methods in Sections III-B and III-C, respectively. As a precursor, Section III-A provides a brief overview of the standard sparse signal recovery problem and the CS framework. Finally, Table I highlights the key distinctions between our proposed CS-based methods and the classical-DFT and MUSIC approaches.
TABLE I: Proposed CS-based methods compared with classical-DFT and MUSIC[12]
Detail Classical-DFT MUSIC Proposed CS-based methods
Radar setup ULA with $\lambda/2$ element spacing ($N_{T,f}$ transmitters and $N_{R,f}$ receivers); $P_{max}$ chirps transmitted for the entire CPI ULA with $\lambda/2$ element spacing ($N_{T,f}$ transmitters and $N_{R,f}$ receivers); $P_{max}$ chirps transmitted for the entire CPI Random SLA with inter-element spacing $\geq\lambda/2$ ($N_{T}<N_{T,f}$ transmitters and $N_{R}<N_{R,f}$ receivers); $P<P_{max}$ chirps transmitted randomly in a CPI
Range estimation DFT with coherent integration of all chirps and array channels Same as classical-DFT DFT with binary integration across all chirps and array channels, or Range-OMP using only one chirp and array channel
Joint Doppler-angle estimation 2D-DFT 2D-MUSIC with spatial smoothing 2D-OMP or vectorized 1D-OMP/ BP/ LASSO
Range resolution $c/2\gamma T_{c}$; increases with total bandwidth swept by LFM chirp Same as classical-DFT Same as classical-DFT for DFT+binary integration; higher for Range-OMP depending on the choice of range grid $\omega_{1\leq g\leq G_{R}}$
Doppler & angular resolutions Doppler resolution ($\lambda/2P_{max}T_{c}$) increases with transmitted chirps; Angular resolution increases with array aperture Higher than classical-DFT; depends on the array aperture, number of chirps, SNR and search grid density Higher than classical-DFT; depends on the choice of Doppler grid $\rho_{1\leq g\leq G_{D}}$ and angular grid $\phi_{1\leq g\leq G_{\theta}}$
Computational complexity $\mathcal{O}(N_{T,f}N_{R,f}P_{max}N\log(N))$ for range estimation; $\mathcal{O}(N_{T,f}N_{R,f}P_{max}\log(N_{T,f}N_{R,f}P_{max}))$ for joint Doppler-angle estimation Same as classical-DFT for range estimation; $\mathcal{O}(N_{T,f}^{2}N_{R,f}^{2}P_{max}+N_{T,f}^{3}N_{R,f}^{3}+G_{1}G_{2}N_{T,f}^{2}N_{R,f}^{2})$ for joint Doppler-angle estimation ($G_{1}$ & $G_{2}$ are Doppler and angular search grid sizes, respectively) Lower than classical-DFT and MUSIC because of smaller number of measurements ($N_{T}<N_{T,f}$, $N_{R}<N_{R,f}$, $P<P_{max}$); see Remark 4 and Table II
@@ III-A CS sparse signal recovery
$\textrm{BP:}\;\;\;\textrm{min}_{\mathbf{x}}\;\;\|\mathbf{x}\|_{1}\;\;\textrm{s.t.}\;\;\|\mathbf{y}-\mathbf{Ax}\|_{2}\leq\epsilon,$ (6)
$\textrm{LASSO:}\;\;\;\textrm{min}_{\mathbf{x}}\;\;\|\mathbf{y}-\mathbf{Ax}\|_{2}\;\;\textrm{s.t.}\;\;\|\mathbf{x}\|_{1}\leq\kappa,$ (7)
@@ III-B Range estimation
We first estimate the target ranges from measurements (3) using two different methods: DFT-based focusing and OMP. Fig. 2 illustrates the conventional and proposed range estimation algorithms. The target velocities and AOAs are then jointly estimated for each detected range in the subsequent section.
@@
$Y_{n,m,p}[l]$ $=\sum_{k=1}^{K}\widetilde{a}^{*}_{k}\exp{(j2\pi\Omega^{k}_{D}\zeta_{p})}\exp{(j2\pi\Omega^{k}_{\theta}(\alpha_{n}+\beta_{m}))}$
$\times\frac{1}{\sqrt{N}}\sum_{t=0}^{N-1}\exp{\left(j2\pi\left(\Omega^{k}_{R}-\frac{l}{N}\right)t\right)}+W_{n,m,p}[l],$ (9)
for $0\leq l\leq N-1$ and $W_{n,m,p}[l]=(1/\sqrt{N})\sum_{t=0}^{N-1}w_{n,m,p}[t]\exp{(-j2\pi lt/N)}$.
@@
$|g(x|\overline{x})|\approx\begin{cases}\sqrt{M},&|x-\overline{x}|\leq\pi/M\omega\\ 0,&|x-\overline{x}|>\pi/M\omega\end{cases}.$ (10)
This approximation implies that in the focus zone $|x-\overline{x}|\leq\pi/M\omega$, the $M$ exponents are coherently integrated while the signal outside the focus zone is severely attenuated. This focusing operation was applied across pulses in [27] to reduce the joint delay-Doppler estimation problem to delay-only estimation in a pulsed-Doppler radar. Particularly, target returns from different pulses were combined into a single high-SNR pulse via time-shifting and modulation. Applying approximation (10) with windowing restricted the focus to a narrow Doppler band, within which delay estimation became feasible. In contrast, our formulation in (9) inherently yields a sum of exponentials in the sampled IF signal’s DFT, without any windowing. We now show that applying approximation (10) to (9) directly concentrates target returns from the same range into a single bin, independent of their Doppler and angular frequencies ($\Omega^{k}_{D}$ and $\Omega^{k}_{\theta}$).
@@
Binary integration: So far, we have considered range estimation from a given $p$-th chirp and $(n,m)$-th virtual array channel. Similarly, the range estimates are computed independently for all $P$ chirps and $N_{T}N_{R}$ channels using the radar measurements $\{y_{n,m,p}[t]\}_{1\leq n\leq N_{T},1\leq m\leq N_{R},1\leq p\leq P}$. In binary integration, these detected ranges are then filtered for false alarms and missed detections across all chirps and channels using a majority rule, i.e., only the ranges detected in the majority of measurements are considered valid. Note that binary integration can also be interpreted as a statistical smoothing technique that mitigates the impact of noise. However, unlike traditional smoothing used in super-resolution methods such as MUSIC, the primary objective here is to suppress false alarms in low-SNR scenarios. In contrast, MUSIC employs smoothing to decorrelate coherent sources and achieve high-resolution estimation at moderate to high SNRs. Moreover, conventional smoothing methods typically involve higher computational complexity compared to binary integration.
Figure 2: Graphical illustration of range estimation using (a) classical-DFT, (b) DFT-focusing with binary integration, and (c) Range-OMP methods.
@@
Classical-DFT range processing also involves peak detection in the DFT of the sampled IF signal. However, in classical processing, the measurements from all chirps and array channels are processed coherently to average out the noise term. At low SNRs, the method fails to detect range bins where the noise is significant. Contrarily, in binary integration, a missed range in one measurement can be detected in other measurements, reducing the missed detection probability, as demonstrated numerically in Section V-A. Note that the classical-DFT method utilizes full measurements from an $N_{T,f}\times N_{R,f}$ ULA and $P_{max}$ chirps, whereas our DFT-focusing approach relies on sparse measurements, leading to reduced computational complexity, as detailed in Table I. However, our joint Doppler-angle estimation procedure proposed later can also be trivially applied to coherently processed measurements.
@@
Consider the measurements $y_{1,1,1}[t]$ from the first chirp and the first virtual array channel. Denote $x_{k}\doteq\widetilde{a}^{*}_{k}\exp{(j2\pi\Omega^{k}_{D}\zeta_{p})}\exp{(j2\pi\Omega^{k}_{\theta}(\alpha_{n}+\beta_{m}))}$ and stack the samples $y_{1,1,1}[t]$ for $0\leq t\leq N-1$ in a $N\times 1$ vector $\mathbf{y}_{R}$. Now, define the $K\times 1$ vector $\widetilde{\mathbf{x}}_{R}=[x_{1},\ldots,x_{K}]^{T}$ such that (3) yields
@@
where the $N\times K$ matrix $\widetilde{\mathbf{A}}(\bm{\Omega}_{R})=[\mathbf{a}(\Omega^{1}_{R}),\ldots,\mathbf{a}(\Omega^{K}_{R})]$ with the $k$-th column $\mathbf{a}(\Omega^{k}_{R})\doteq[1,\exp{(j2\pi\Omega^{k}_{R})},\ldots,\exp{(j2\pi\Omega^{k}_{R}(N-1))}]^{T}$. Here, $\mathbf{w}_{R}$ represents the $N\times 1$ stacked noise vector. Each column $\mathbf{a}(\Omega_{R})$ of matrix $\widetilde{\mathbf{A}}$ is parameterized by $\Omega_{R}$ and is the equivalent steering vector in the (range) beat frequency domain. Furthermore, $\mathbf{y}_{R}$ in (13) represents full measurements in the fast-time domain sampled at frequency $f_{s}$.
@@
where the $G_{R}\times 1$ vector $\mathbf{x}_{R}$ contains the target ranges and unknown coefficients $\{x_{k}\}$. In particular, a non-zero element of $\mathbf{x}_{R}$ represents a target present at the range corresponding to the grid point. Since $K\ll G_{R}$, $\mathbf{x}_{R}$ is a sparse vector and range estimation reduces to determining $\textrm{supp}(\mathbf{x}_{R})$ given $\mathbf{y}_{R}$ and $\mathbf{A}$. To this end, we can employ any 1D-OMP or BP recovery algorithms described in Section III-A. In particular, we use OMP in Section V. Note that our Range-OMP employs the standard OMP algorithm but utilizes full measurements in the radar’s fast-time domain, rather than compressed ones. It is used here to achieve high-resolution range estimation with reduced computational cost compared to conventional DFT-based methods.
@@
Range-OMP in (14) considers measurements only from one chirp and one array channel. This is in contrast to DFT-focusing, wherein all the measurements are integrated (binary or coherently) to obtain valid target ranges making them computationally expensive. Furthermore, Range-OMP provides superior range resolution compared to DFT-focusing, determined by the choice of grid $\omega_{1\leq g\leq G_{R}}$ and not by parameters $\gamma$ and $T_{c}$. However, the $\textrm{supp}(\mathbf{x}_{R})$’s recovery still depends on the matrix $\mathbf{A}$ such that the recovery probability is low if the grid points are too close. Note that while binary integration can also be employed after Range-OMP to suppress false targets at low SNRs, this step becomes unnecessary in high-SNR scenarios.
Remark 4 (Computational complexity).
The computational complexity of DFT-focusing depends on the number of DFT points $N$, the number of virtual array channels $N_{T}N_{R}$, and the number of pulses $P$ per CPI. Specifically, computing the complex DFT has a complexity of $\mathcal{O}(N_{T}N_{R}PN\log(N))$[50], followed by binary integration with complexity $\mathcal{O}(N_{T}N_{R}PN)$[10]. Although this matches the complexity order of conventional DFT-based range estimation using full measurements (Table I), our DFT-focusing approach incurs a lower computational burden due to the significantly reduced number of measurements $N_{T}N_{R}P$, enabled by the random SLA and sparse chirps. In contrast, our Range-OMP uses measurements from only a single channel and chirp, and follows the standard OMP framework. Its complexity is determined by the number of targets $K$, the size of the range grid $G_{R}$, and the number of DFT points $N$, resulting in a complexity of $\mathcal{O}(KNG_{R})$[51]. Moreover, using a finer range grid (larger $G_{R}$) enables better range resolution while incurring lower computational cost than DFT-based methods, as the number of targets $K$ is much smaller than the number of measurements $N_{T}N_{R}P$.
@@ III-C Joint Doppler-angle estimation
We stack the measurements $Y_{n,m,p}[l^{\prime}]$ for all $1\leq n\leq N_{T}$ and $1\leq m\leq N_{R}$ in an $N_{T}N_{R}\times 1$ vector $\mathbf{y}_{p}$ for the $p$-th chirp. Further, define the $N_{T}N_{R}\times P$ matrix $\mathbf{Y}=[\mathbf{y}_{1},\ldots,\mathbf{y}_{P}]$ and the steering vectors in the Doppler and angular frequency domains as $\mathbf{b}(\Omega_{D})$ and $\mathbf{c}(\Omega_{\theta})$, respectively, given by
@@
where the unknown $G_{\theta}\times G_{D}$ matrix $\mathbf{Z}$ contains the coefficients $\{z_{k^{\prime}}\}$ as well as the target velocities and AOAs information. Particularly, a non-zero element in $\mathbf{Z}$ represents a target with its AOA and velocity equal to that of the corresponding point in the 2D grid $\{(\phi_{i_{1}},\rho_{i_{2}}):1\leq i_{1}\leq G_{\theta},1\leq i_{2}\leq G_{D}\}$. Since $K^{\prime}\ll G_{D}G_{\theta}$, the matrix $\mathbf{Z}$ is sparse, and hence, the desired velocities and AOAs can be recovered using small number of measurements $\mathbf{Y}$ given matrices $\mathbf{C}$ and $\mathbf{B}$. Note that the joint Doppler-angle estimation has now been reduced to determining $\textrm{supp}(\mathbf{Z})$. In (19), the measurement matrices $\mathbf{C}$ and $\mathbf{B}$ and hence, the recovery guarantees, depend on the choice of girds $\rho_{1\leq g\leq G_{D}}$ and $\phi_{1\leq g\leq G_{\theta}}$ as well as the number and (random) locations of the transmitted chirps ($\zeta_{1\leq p\leq P}$) and antenna elements ($\alpha_{1\leq n\leq N_{T}}$ and $\beta_{1\leq m\leq N_{R}}$).
TABLE II: Computational complexity of joint Doppler-angle estimation ($N_{\text{int}}$ is the number of iterations)
Method Complexity Core Operation
Vectorized 1D-OMP $\mathcal{O}(KN_{T}N_{R}PG_{\theta}G_{D})$ Projection and least-squares
Vectorized BP $\mathcal{O}(G_{\theta}^{3}G_{D}^{3})$ Linear programming (e.g., interior point)
Vectorized LASSO $\mathcal{O}(N_{\text{int}}N_{T}N_{R}PG_{\theta}G_{D})$ Convex optimization (e.g., ISTA, FISTA)
2D-OMP $\mathcal{O}(K(N_{T}N_{R}+P)G_{\theta}G_{D})$ Matrix projection and least-squares
@@
which is same as (4) with $\mathbf{A}$ and $\mathbf{x}$ replaced by $\mathbf{D}$ and $\mathbf{z}$, respectively. In Section IV, we consider the $l_{1}$-minimization-based solution of (20) to derive the theoretical recovery guarantees while Section V numerically compares 2D-OMP and vectorized 1D-OMP, BP, and LASSO methods. Notably, with fixed Doppler and angular grids, these methods exhibit similar detection performance and estimation accuracy. However, their computational complexities vary, as summarized in Table II. The complexity of all approaches primarily depends on the number of grid points $G_{D}$ and $G_{\theta}$, the number of virtual array channels $N_{T}N_{R}$, the number of chirps $P$, and the number of targets $K$, but the vectorized BP method incurs the highest computational cost. This is because $G_{D}G_{\theta}$ is typically set to a large value to achieve fine Doppler and angular resolution. In contrast, 2D-OMP achieves a computational advantage through direct matrix operations, offering a speedup factor of $N_{T}N_{R}P/(N_{T}N_{R}+P)$ over vectorized 1D-OMP[48].
@@ IV-A Uniform recovery
Then, for $0<\vartheta<1$, the coherence $\mu$ of matrix $\mathbf{D}$ in (20) satisfies the following:
@@
$\times(1-2\vartheta\sqrt{N_{T}N_{R}}\;\mathcal{K}_{1}(2\vartheta\sqrt{N_{T}N_{R}}))^{G_{\theta}-1},$ (28)
@@
Consider the random sparse chirps and SLA radar setup with the distributions $\mathcal{P}_{\alpha}$, $\mathcal{P}_{\beta}$ and $\mathcal{P}_{p}$ and grids $\rho_{1:G_{D}}$ and $\phi_{1:G_{\theta}}$ satisfying the conditions of Theorem 1. Also, consider some $0<\epsilon<1$ with two positive constants $\epsilon_{1}$ and $\epsilon_{2}$ such that $\epsilon_{1}+\epsilon_{2}=\epsilon$ and
@@
$N_{T}N_{R}\geq\kappa_{2}\left(K-\frac{1}{2}\right)^{2}$
$\;\;\;\times\left(\log\left(\frac{G_{\theta}\sqrt{\pi}}{2\epsilon_{2}}\right)+\frac{1}{2}\log\left(2\log\left(\frac{G_{\theta}\sqrt{\pi}}{2\epsilon_{2}}\right)\right)\right)^{2},$ (31)